Find f(y) so that equation f(y)dx − zx dy − xy ln y dz = 0 is integrable. Also obtain the corresponding integral using Natani’s method.
1
Expert's answer
2020-03-16T13:56:13-0400

Equation "f(y)dx-zxdy-xylnydz=0" is integrable, if 

"X\cdot rotX=0\\
X=(f(y),-zx,-xylny)"

Calculate "rotX" :

"rotX=\nabla\times X=\begin{vmatrix}
 \vec{i} & \vec{j}&\vec{k} \\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y}& \frac{\partial}{\partial z}\\
f(y)&-zx&-xylny
\end{vmatrix}=\\
=\vec{i}( \frac{\partial}{\partial y}(-xylny)- \frac{\partial}{\partial z}(-xz))-\\
-\vec{j}( \frac{\partial}{\partial x}(-xylny)- \frac{\partial}{\partial z}(f(y)))+\\
+\vec{k}( \frac{\partial}{\partial x}(-zx)- \frac{\partial}{\partial y}(f(y)))=\\
=(-xlny)\vec{i}+(ylny)\vec{j}+(-z-f'(y))\vec{k}"

Then

"X\cdot rotX=-xlny\cdot f(y)+xylnyf'(y)=0"

Thus "xlny\cdot(yf'(y)-f(y))=0" , hence "yf'(y)-f(y)=0"

The solution of the last equation is determined by

"\frac{df}{f}=\frac{dy}{y}\\
ln|f(y)|=ln|y|+ln|A|\\
f(y)=A\cdot y"

where "A" is a real constant. Finally, equation takes the form 

"Aydx-zxdy-xylnydz=0"

Using Natani’s method we firstly put "z=const" and solve the obtained equation

"Aydx=zxdy\\
a\frac{dx}{x}=z\frac{dy}{y}\\
Alnx=zlny+F(z)"

Now we let "x=1" . Then equation takes the form 

"-zdy-ylnydz=0"

and

"F(z)=-zlny"

Thus

"zdy+ylnydz=0\\
\frac{dy}{ylny}+\frac{dz}{z}=0\\
ln|lny|=-ln|z|+ln|C|\\
lny=\frac{C}{z}"

Therefore, 

"F(z)=-zlny=-z\cdot\frac{C}{z}=-C=K=const"

Finally,

"Alnx=zlny+K"

where "K" is an arbitrary real constant.